the flagpole is 18 ft high and a model of the same flag pole is 2 in high. what is the ratio of the model to the real height?
Answer by TimothyLamb(4379)

2/(18*12) = 1/(18*6) = 1/108
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---